OpenStudy (mathmath333):

Is their a formula for finding finite sum of series with changing differences ?

OpenStudy (mathmath333):

\[Ex\] \[1+3+6+10+15+21.....\]

OpenStudy (mathmath333):

\[3+10+19+30+43...\]

Nnesha (nnesha):

there is two formula for finite series one for arithmetic and one for geometric

OpenStudy (mathmath333):

i know that but this series is different ,it shows changing difference pattern

Nnesha (nnesha):

yeah i mean there is a pattern but not same difference or same ratio but they have a pattern is that question you have or you just make that

OpenStudy (mathmath333):

is their formula for finding the nth term here if its a sequence

OpenStudy (mathmath333):

if yes then their should be formula for finding summation too

Nnesha (nnesha):

i don't think so we just have to figure out

Nnesha (nnesha):

for example this one 5,10,17,26,3 pattern for this one 5+5,10+7,17+11 mean we adding odd numbers every time

OpenStudy (mathmath333):

this is too complex for me

OpenStudy (xapproachesinfinity):

You can recognize that we are adding asequence which known to us 1,2,3,4,......n

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

@geerky42

OpenStudy (xapproachesinfinity):

observe that \(1+3+6+10+15+21.....=1+(1+2)+((1+2)+3)+(((1+2)+3)+4)+((((1+2)+3)+4)+5).....\)

OpenStudy (xapproachesinfinity):

........ (forgot the continuation lol)

OpenStudy (xapproachesinfinity):

we can see that after taking care of parenthesis we infinite sum of 1+1+1+1.... 2+2+2+2+2.... 3+3+3+3+3 and so on

OpenStudy (xapproachesinfinity):

n sums of 1 is just n n sums of 2 is 2n n sums of 3 is 3n and so on then we would have n+2n+3n+4n+5n+6n........+n^2...... if we factor out n we would have n(1+2+3+4+5+6.....+n) for first n! and we already know the sum of \(\sum_{i=1}^{n}i=\frac{n(n+1)}{2}\) so the sum of that series for the first n numbers is \(n\sum_{i=1}^{n}i=\frac{n^2(n+1)}{2}\)

OpenStudy (xapproachesinfinity):

i guess this is how you would look at it! but not sure if i'm doing it the right way hehe

OpenStudy (xapproachesinfinity):

let's test this if it is true hehe

OpenStudy (xapproachesinfinity):

does not go right hehe

OpenStudy (mathmath333):

\(\large\ =\dfrac{n(n+1)(n+2)}{6}\)

OpenStudy (xapproachesinfinity):

where did you come with that that's for power series \(\sum_{k=1}^{n}k^2\)

OpenStudy (xapproachesinfinity):

this a fun series! 1+1+1+1+....+2+2+2+2+.....+3+3+3+3+..... what could be the sum of this! i was confident that i was going in the right path hehehe

OpenStudy (mathmath333):

\(\large\bf u~~~ can~~~ check~~~ it\) \(\large\sum _{i=1}^{\infty}\ n^2=\frac{n(n+1)(2n+1)}{6}\)

OpenStudy (xapproachesinfinity):

well it is blowing to infinity, but can we have a finite sum for a finite numbers!

OpenStudy (mathmath333):

yes

OpenStudy (xapproachesinfinity):

No that's a different case we have the other series 1+2+3+4... not \(1^2+2^2+3^2+4^2.....\)

OpenStudy (mathmath333):

but good to see u tried hard....

OpenStudy (xapproachesinfinity):

tag some other people hehe

OpenStudy (mathmath333):

no need

OpenStudy (mathmath333):

at last got it

OpenStudy (xapproachesinfinity):

hmmm, you sure! i think there are people who can help much better and this stuff are fun to think about

OpenStudy (xapproachesinfinity):

@ganeshie8 , @amistre64

OpenStudy (mathmath333):

i know they can solve in seconds

OpenStudy (amistre64):

yes there is lol

OpenStudy (xapproachesinfinity):

well you question is kinda general!

OpenStudy (amistre64):

1+3+6+10+15+21 find the first differences (akin to a first derivative if you will) 1+3+6+10+15+21 2 3 4 5 6 now the second differences are constant, a constant second derivative leads us to a quadratic right?

OpenStudy (mathmath333):

yes

OpenStudy (xapproachesinfinity):

riemann sums?

OpenStudy (amistre64):

we can work this by taking the leading values (1,2,1) and applying them as: 1/0! + 2n/1! + 1[n(n-1)]/2! and simplify as needed, this is for n>=0

OpenStudy (amistre64):

its akin to creating a taylor series

OpenStudy (xapproachesinfinity):

oh i see!

OpenStudy (xapproachesinfinity):

now the second differences are constant, a constant second derivative leads us to a quadratic right? ========= i didn't understand your line here?

OpenStudy (mathmath333):

\(\large\ n^2-n+1\)

OpenStudy (amistre64):

a discrete formula has rates of change that are called differences a continuous rate of change are called derivatives

OpenStudy (mathmath333):

its for nth term

OpenStudy (amistre64):

you might have to adjust it if your n starts at 1

OpenStudy (amistre64):

if n starts at 1, and we are defined for n starts at 0, then let k=n-1, n=k+1

ganeshie8 (ganeshie8):

mathmath333 you should also think about triangular numbers they did not give you just some random quadratic sequence, the sequence is a nice well known triangular numbers

OpenStudy (xapproachesinfinity):

and how did you decide that is quadratic!

OpenStudy (mathmath333):

@ganeshie8 i absolutely now this

ganeshie8 (ganeshie8):

second sequence looks bit off though

OpenStudy (xapproachesinfinity):

eh! i'm lost now! hehe

OpenStudy (mathmath333):

but its right

OpenStudy (amistre64):

f = x^2 f' = 2x f'' =2 notice that a quadratic has a constant 2nd derivative

OpenStudy (xapproachesinfinity):

yes! i see that

OpenStudy (amistre64):

a constant 2nd differences will be a discrete quadratic

OpenStudy (mathmath333):

@ganeshie8 if i know the formula for nth term for 2nd series then can i find the summaation of the second series?

OpenStudy (amistre64):

we can work a long way: a0 a1 a2 a3 a4 b0 b1 b2 b3 c c c bn = b0 + nc a0 a1 a2 a3 a4 b0+0c b0+1c b0+2c b0+3c a1 = a0 + b0 + 0c a2 = a0 + b0 + 0c + b0+1c = a0 + 2b0 + c(0+1) a3 = a0 + 3b0 + c(0+1+2) a4 = a0 + 4b0 + c(0+1+2+3) an = a0 + nb0 + c(1+2+3+n-1)

OpenStudy (amistre64):

forgot a .... inthe c parts lol

ganeshie8 (ganeshie8):

yes you can find partial sum for any sequence : \[\large \sum \limits_{n=1}^{k} a_n\] whats your formula for nth term ?

OpenStudy (xapproachesinfinity):

hmm now it looks way better! thanks @amistre64

OpenStudy (amistre64):

if you work enough of them you see a correlation to differential equations

OpenStudy (amistre64):

the difference tree also creates a kind of pascal triangle

ganeshie8 (ganeshie8):

thats like finding 3 equations and solving the unknowns : 1 3 6 10 let the required nth term formula be "ax^2+bx+c" you can get 3 equations by using the first 3 terms : a + b + c = 1 4a + 2b + c = 3 9a + 3b + c = 6 we can solve these, but yes you won't learn anything about the pattern with this method

OpenStudy (mathmath333):

\(\large\ nth~~term=n^2+4n-2\)

OpenStudy (xapproachesinfinity):

oh right! that looks like a pascal triangle!

OpenStudy (xapproachesinfinity):

i had less practice on differential equations! hehehe

ganeshie8 (ganeshie8):

partial sum of first \(k \) terms \[\large \sum \limits_{n=1}^{k} (n^2 + 4n-2) = \sum \limits_{n=1}^{k}n^2 + 4\sum \limits_{n=1}^{k} n -2\sum \limits_{n=1}^{k} 1 \]

ganeshie8 (ganeshie8):

you must be having readymade formulas for sum of squares and sum of first n terms already ?

OpenStudy (mathmath333):

yes i know few

ganeshie8 (ganeshie8):

its time to use them to your advantage

OpenStudy (amistre64):

if not, you can always use what we did at the start to develop them :)

ganeshie8 (ganeshie8):

\[\large \begin{align} \sum \limits_{n=1}^{k} (n^2 + 4n-2) &= \sum \limits_{n=1}^{k}n^2 + 4\sum \limits_{n=1}^{k} n -2\sum \limits_{n=1}^{k} 1 \\~\\ &= ? + ? + ? \end{align} \]

ganeshie8 (ganeshie8):

see if you can simplify and get a nice looking expression for the sum in terms of "k" alone

OpenStudy (xapproachesinfinity):

eh he already mentioned the two sum N^2 and N the sums of 1's are easy

OpenStudy (xapproachesinfinity):

@mathmath333, just check you replies above^_^

OpenStudy (mathmath333):

lol i will

ganeshie8 (ganeshie8):

here in one place : \(\large \sum \limits_{n=1}^{k}n^2 = \dfrac{k(k+1)(2k+1)}{6} \) \(\large \sum \limits_{n=1}^{k}n = \dfrac{k(k+1)}{2} \) \(\large \sum \limits_{n=1}^{k}1 = k \)

OpenStudy (mathmath333):

thanks everyone

OpenStudy (mathmath333):

oh its wrong hehe

ganeshie8 (ganeshie8):

whats wrong ?

OpenStudy (mathmath333):

nothing , i posted a wrong equation

OpenStudy (mathmath333):

\(\large\bf\ =\dfrac{k(2k^2+15k+1)}{6}\)

OpenStudy (mathmath333):

at last........sigh

OpenStudy (mathmath333):

very happy